Tylan oxide update

maurice maurice at stanford.edu
Fri Jan 7 17:20:56 PST 2011


Tylan1-  controller has been sent to tystar for repair...should be back 
late next week.
-
Tylan2- controller returned from vendor today and was installed.  During 
the first test there was an O2Hi error.  Troubleshooting will continue 
on Monday
-
Tylan3- Up
-
Tylan4- Wet ox is thin...System will be yellow over the weekend.



More information about the tylan1 mailing list